### v1

# Graph Theory and GNNs to Unravel the Topographical Organization of Brain Lesions in Variants of Alzheimer’s Disease Progression

Abstract
This study utilizes graph theory and deep learning to assess variations in Alzheimer’s disease (AD) neuropathologies, focusing on classic (cAD) and rapid (rpAD) progression forms. It analyses the distribution of amyloid plaques and tau tangles in postmortem brain tissues. Histopathological images are converted into tau-pathology-based graphs, and derived metrics are used for statistical analysis and in machine learning classifiers. These classifiers incorporate SHAP value explainability to differentiate between cAD and rpAD. Graph neural networks (GNNs) demonstrate greater efficiency than traditional CNN methods in analyzing this data, preserving spatial pathology context. Additionally, GNNs provide significant insights through explainable AI techniques. The analysis shows denser networks in rpAD and a distinctive impact on brain cortical layers: rpAD predominantly affects middle layers, whereas cAD influences both superficial and deep layers of the same cortical regions. These results suggest a unique neuropathological network organization for each AD variant.
Index Terms—  Alzheimer, Graph Neural Network, XAI
1 Introduction
Alzheimer’s disease (AD), a leading neurodegenerative disorder, is characterized by memory loss, cognitive decline, and behavioral changes, mainly in the elderly. The disease is believed to be caused by the accumulation of beta-amyloid proteins forming amyloid plaques and altered tau proteins leading to neurofibrillary tangles, disrupting neuronal communication. AD’s complexity stems from the brain’s intricate structure and the disease’s heterogeneity, with patients experiencing either a rapid, aggressive pathology (rpAD) or a slower classical decline (cAD). Diagnosing AD, often similar to senile dementia, is typically confirmed postmortem through histopathological examination, identifying amyloid plaques and tau tangles in brain tissue, a method limited by its postmortem nature, high costs, and potential for inconsistencies in manual annotation.
Whole Slide Images (WSI) are high-resolution scans crucial in Alzheimer’s disease research for examining the cerebral cortex’s six distinct layers. Despite AI’s success in disease pattern analysis, the large size of WSIs poses processing challenges, which can be addressed through innovative graph representation. This approach significantly reduces data size while preserving spatial integrity and offers an unbiased, orientation-agnostic representation. Graph topological metrics help analyze tissue structures, providing insights into disease characteristics as noted in [1].
Graph Neural Networks (GNNs) have emerged as critical tools in bioinformatics and neurobiology, particularly for their effectiveness with graph-structured data, key in interpreting biological networks and neurodegenerative diseases [2, 3]. In AD research, GNNs have shown promise in patient classification and staging, with applications like the multi-class Graph Convolutional Neural Network (GCNN) for categorizing disease progression [4, 5, 6, 7]. Their growing significance in AD detection and prognosis is also evident [8], alongside their role in patient stratification and the advancement of individualized treatments [9, 10, 9]. Further reports on the performance and potential of various GNN architectures have contributed significantly to AD research [11, 12, 13], with applications also in tissue abnormality characterization [14, 15, 16], enhancing our understanding of AD’s pathology.
The primary objective of this study is to discern the underlying distinctions between rapid and classic Alzheimer’s disease progression by focusing on the distribution patterns of plaques and tangles across different brain tissue layers. Moving beyond the morphological characteristics of these elements [1], we emphasize their topographical arrangement within the tissue. This approach offers novel insights into the physiopathology of AD’s onset and progression, potentially leading to improved diagnostic methods and targeted therapeutic interventions (precision medicine).
2 Methods
2.1 Statistical Analysis & Machine Learning
This study involves the creation of two graph sets: one representing patient-level data and the other at the layer level or subgraphs. Two clustering methods—Connected Components and Markov Clustering—were applied for the patient-level set. Both plaque and tangle objects were independently represented in these graphs and subgraphs. Various metrics, such as node and edge counts, total edge length, cluster characteristics, and $\alpha_{optimal}$ values, were computed from these structures. These metrics, normalized relative to the Region of Interest (ROI) size (except for the raw and normalized $\alpha$ values), provided insights into the structural and spatial dynamics within WSI. Additionally, each metric was normalized to the $\alpha_{optimal}$ value, ensuring consistent analysis across varying WSI and accounting for graph size and ROI differences.
These metrics were then used as features in a Random Forest (RF) classifier to distinguish between cAD and rpAD. The classifier’s accuracy was enhanced by using recursive feature elimination, analyzing feature importance, and applying SHAP values for in-depth data and model decision interpretation. We also meticulously addressed potential data leakage by recognizing and handling dependencies among clusters from the same graph.
In addition, statistical analysis employed distribution plots, correlation matrices, and multivariate techniques to discern subtle variations between cAD and rpAD. This approach aimed to deepen the understanding of the disease’s patterns.
2.2 Deep Learning Methods
This study utilizes Deep Learning to create meaningful clusters with node embeddings from Graph Neural Network (GNN) latent spaces, employing explainable AI (XAI) to discern model-relevant structures. A variational graph auto-encoder (VGAE) was deemed unsuitable due to the artificial nature of graph construction, leading to the selection of a GNN classifier.
A. GNN Classifier:
The focus was on the spatial arrangement of plaques and tangles within the WSI, independent of their morphological attributes. Graph theoretical measures generated node features such as degree, clustering coefficient, betweenness, closeness, and layer location, with edge weights inversely proportional to distance. For training our GNNs, we used PyTorch Geometric and k-fold cross-validation to ensure our results were reliable. Various GNN architectures were tested, including Graph Attention Network (GAT), Graph Convolution Network (GCN), Graph Sample and Aggregation (GraphSAGE), Weighted GraphSAGE, and Chebyshev Spectral (ChebNet). The models provided node embeddings or classifications with an early stopping mechanism to avoid overfitting. Optimal parameters were determined through hyperparameter tuning. Model efficacy was measured using a weighted score of accuracy and loss. Figure 1 summarizes the models implemented.
Fig. 1: Model architecture and different output options.
B. Clustering on Node Embeddings
We used the 12-dimensional node embeddings extracted from the final convolution layer of our GNN for clustering purposes. These embeddings are complex representations of the nodes, reflecting the features and interrelations the GNN learned during its training. The precision with which each node’s role within the network is captured is reflected in the model’s accuracy. Clusters derived from these node embeddings offer greater significance and reliability than those formed directly from the graph data. This approach notably augments the Random Forest Classifier’s precision and depth of analysis, particularly in distinguishing between cAD and rpAD cases.
C. GNN Explainer
GNNExplainer and PGExplainer are applied post-training to analyze the model’s decision-making process. The GNNExplainer offered instance-level explanations identifying the nodes and edges relevant to the prediction. On the other hand, PGExplainer provided a broader probabilistic perspective by identifying relevant subgraphs within the input graph that contribute to the model’s predictions. This dual analysis served as a form of cross-validation, building trust in the validity of our findings.
3 Dataset and Graph Construction
This study used a dataset of semi-automatically annotated WSI, featuring samples from 18 patients—12 diagnosed with cAD and 6 with rpAD. These images were sourced from the French National Brain Biobank Neuro-CEB, with requisite consent obtained from patients or their next of kin. Frontal lobe sections, stained using AT8 antibody to accentuate phosphorylated tau, were digitized via Hamamatsu scanners at resolutions of 227 nm/pixel and 221 nm/pixel. The WSIs display a spectrum of AD pathology, varying in tau pathology severity, staining quality, and tissue preservation. The dataset also includes detailed features like the object’s coordinates, area, size of plaques and tangles, and the layers to which they belong. Notably, the dataset presents a wide variance in both the size of the WSIs and the count of tangles and plaques, ranging from dozens to thousands. Despite its small size, the dataset’s high quality and expert annotations form a solid base for our analysis. Figure 2 illustrates sample annotations.
Graphs were generated focusing on the spatial relationships of plaques and tangles, excluding morphological characteristics. Delaunay triangulation was employed, connecting points to their nearest neighbors, thereby minimizing hyperparameter selection bias. This approach is ideal for biological studies due to its natural adaptation to point density and resilience against noise and outliers.
We applied erosion to the graphs to isolate core structures and components within the graph. The erosion is guided by an optimal alpha value derived from the data, controlling the level of granularity in the graph representation. This alpha value, set at half the median of all values from the Delaunay graph ($\alpha_{optimal}=\frac{1}{2}\cdot median(\alpha)$), ensured consistent erosion, emphasizing biologically pertinent structures and omitting less significant elements [17]. Edges exceeding 1000 micrometers were excluded to align with the biological understanding that cellular interactions beyond this range are improbable.
Edge weighting in the graphs was inversely proportional to the square root of the distance between nodes, emphasizing the significance of proximity in cellular interactions. This approach in non-directed graphs reflects the bidirectional nature of these interactions, offering a more authentic representation of the biological networks under study.
To counter the imbalance between cAD and rpAD cases in the dataset, rpAD cases were oversampled. This strategy maintained balance in model training without altering the intrinsic properties of the data, avoiding the pitfalls of conventional data augmentation methods.
Fig. 2: Example of WSI images and its annotations on the Global Region of Interest (ROI) on the left and the 6 layers delimitation on the right.
4 Results
Our analysis, integrating statistical methods and machine learning classifiers, successfully differentiated between cAD and rpAD. The cluster data visualization indicated notable variance in feature distributions and correlations between the two conditions. Using different clustering algorithms, the RF classifier attained accuracies between 70
In evaluating five GNN models, the GCN, GraphSAGE, and ChebNet models demonstrated robust performance. In contrast, the GAT faced challenges with limited node features, and the weighted version of GraphSAGE produced inconsistent outcomes. GCN, GraphSAGE, and ChebNet were selected for in-depth analysis due to their varied latent space representations.
Given the dataset’s limited size, the models exhibited variability in performance across runs. To mitigate this, each model was executed multiple times, selecting node embeddings from the epoch with the highest performance metrics in the most accurate model. This process highlights the approach’s potential while acknowledging inherent variability that may impact reliability and generalizability.
The RF classifier, enhanced with node embeddings clustering, achieved an impressive accuracy range of 95-99
Fig. 3: SHAP values visualization of the RFC differentiating between cAD and rpAD plaque clusters derived from GNN node embeddings. A positive impact pushes the model towards predicting rpAD, and inversely a negative impact to cAD.
The congruence in node and edge importance metrics, as identified by both GNN Explainer and PGExplainer, adds validity to our findings, suggesting they accurately represent underlying patterns. As shown in Figure 4, the importance distribution across different brain layers corroborated insights from the RF classifier, indicating that intermediate layers hold crucial discriminative information. Furthermore, the analysis revealed distinct importance levels across layers, with outer layers (2, 5, and 6) being more significant for cAD and inner layers (3 and 4) for rpAD. This observation implies unique structural or informational characteristics pertinent to each Alzheimer’s Disease subtype at various brain depths.
Fig. 4: Mean node and edges importance for cAD and rpAD for GNNExplainer.
5 Conclusion and Discussion
This research offers significant insights into AD, particularly in differentiating between cAD and rpAD. One key observation is the denser and more interconnected graph structures in rpAD compared to cAD. Specifically, rpAD exhibits compact clusters as the number of objects increases, contrasting with the expanding clusters in cAD. Additionally, we observed distinct effects on brain layers: rpAD predominantly affects middle layers, whereas cAD primarily influences outer layers. These findings imply divergent cognitive impacts and clinical presentations between the two AD subtypes, mirroring each variant’s unique neuropathology.
GNN methodologies proved advantageous over traditional analysis techniques, offering enriched, reliable insights into complex disease patterns. While traditional methods provide a foundational comparison, GNNs facilitate a more profound understanding of AD’s intricate dynamics. For instance, GNN embeddings, as depicted in Figure 5, present a more coherent spatial distribution, aligning with the brain’s layered structure, in contrast to the arbitrary groupings typical of classical clustering.
Fig. 5: Comparison of clustering method for raw data points and GNN-based embeddings. The top graph presents clustered nodes based on their spatial coordinates. In contrast, the bottom graph leverages GNN embeddings to cluster nodes.
Looking forward, integrating Voronoi diagrams (complementary to Delaunay triangulation) could enhance spatial analysis. Enlarging the dataset would refine differentiation and bolster model reliability. Furthermore, incorporating three-dimensional (3D) data to reconstruct brain regions and develop 3D graphs could unveil connections obscured in two-dimensional representations, yielding more comprehensive insights. Such advancements accentuate GNNs’ ability to capture spatial relationships in multi-dimensional spaces, potentially offering perspectives not fully captured by CNNs.
In conclusion, our approach combined statistical precision, machine learning’s predictive capabilities, and deep learning’s analytical depth to explore the nuances of AD progression. By employing a multi-faceted methodology, we aimed to ensure a thorough, wide-ranging analysis capable of discerning the subtle variances between cAD and rpAD in brain tissue. Although constrained by the dataset’s size, our findings and methodologies lay the groundwork for more expansive studies, emphasizing GNNs’ potential in elucidating complex neurological disorders.
6 Acknowledgments
This research was supported by Mr Jean-Paul Baudecroux and The Big Brain Theory Program - Paris Brain Institute (ICM). The human samples were obtained from the Neuro-CEB brain bank (https://www.neuroceb.org/en/) (BRIF Number 0033-00011), partly funded by the patients’ associations ARSEP, ARSLA, “Connaître les Syndromes Cérébelleux”, France-DFT, France Parkinson and by Vaincre Alzheimer Fondation, to which we express our gratitude. We are also grateful to the patients and their families.